Prev Next

SW&IT Meeting Minutes 2021-03-08

Attendees: David, Igor, Matthieu, Michael, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210222
LR20210222 - Bec, Matthieu 64.0
LR20210222 - Edmeades, David 26.0
LR20210222 - Sola, Igor 0.0
LR20210222 - Grenz, Paul 4.0
LR20210222 - Hooper, Stephen 80.0
LR20210222 - Zhang, Yang 81.0
LR20210222 - Zhang, Xianyu 80.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • Igor handover
    • ATT tasks and priorities
    • SW student induction
  • This week
    • ATT PM priorities, MCS, candidates review/interview
    • Igor handover ctd.
    • dms refactoring, websockets
    • $ADOPT_LOG examination
    • OT SHARK-NIR

XZ

  • Last week
    • Analyze SX AdSec issues
    • ATT AdSec and SW meeting
    • Install and configure SOUL msgd fix in SX AdSec
    • AO support
  • This week
    • SX AdSec issues/test
    • SOUL test support
    • AO support

YZ

  • Previous:
    • OSCO git/jenkins work (it's working now!)
    • Web-DMS dev.
    • OT/PIT meeting with Igor and other SWG/Sciops members
    • IT#8329 (RB_Science issue reported by Jenny). Cannot reproduce, multiple tests and its okay. So maybe a user error.
    • Some MODS and LBC tweaks for Olga.
    • TMS meeting and some planning.
    • IT/Sciops robs roundup meeting Tuesday.
  • Next:
    • Web-DMS dev (to continue)
    • OT/PIT/LUCI meeting.
    • Possible OSCO meeting.
    • TMS test and TCS operations review with Heejoo on Monday.
    • TMS extraction utility upgrade.
    • TMS software setting tweak and wiki updates.
    • RB_Science tweaks (if have time).
    • TCS/AOS patch release.
    • Sick day on Friday.

DE

  • Last week:
    • Synology work
      • eBay unit picked up by UPS, refund issued from seller, not our problem anymore
      • RMA unit seems to be impossible for an end user to install anything but latest firmware, so we are forced into a downtime on the production cluster, as discussed in ITmeeting. Time pending. This is apparently uncharted territory for Synology support.
    • Storage focused IT meeting
      • Defined tasks to gather a current picture of proposed data burn rate for new instruments
      • Offsite backup options
    • Sales calls with Dell and Backblaze
    • Laura BT KB issue. Again. Solved for the moment with USB, but I put in a call to my new, eager Dell contact to jump up the support queue.
    • Autocad installs for Robert student
  • This week:
    • Final reinstall of DSM on RMA Synology unit in prep for switchover
    • Much documentation for tasks mentioned above
    • Received rack console, will check it out for installation when I go in for the Synology
    • Mac Mini status?
    • Cobbler

SH

  • DIMM hardware (no progress)
  • network config into git
  • VM stuff (little progresS)
  • robs meeting
  • Igor meetings
  • IT meeting

next:

  • Mac IP and DNS (from IT meeting)
  • laptop for Mike
  • begin phase II move to LDAP
Topic revision: r11 - 08 Mar 2021, YangZhang
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback